# Inference-Time Text-to-Video Alignment with Diffusion Latent Beam Search

### v1

@@ Abstract
The remarkable progress in text-to-video diffusion models enables photorealistic generations, although the contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes. Recently, an alignment problem has attracted huge attention, where we steer the output of diffusion models based on some quantity on the goodness of the content. Because there is a large room for improvement of perceptual quality along the frame direction, we should address which metrics we should optimize and how we can optimize them in the video generation. In this paper, we propose diffusion latent beam search with lookahead estimator, which can select better diffusion latent to maximize a given alignment reward, at inference time. We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration by weighting existing metrics. When evaluating outputs by using vision language models as a proxy of humans, many previous metrics to quantify the naturalness of video do not always correlate with evaluation and also depend on the degree of dynamic descriptions in evaluation prompts. We demonstrate that our method improves the perceptual quality based on the calibrated reward, without model parameter update, and outputs the best generation compared to greedy search and best-of-N sampling. We provide practical guidelines on which axes, among search budget, lookahead steps for reward estimate, and denoising steps, in the reverse diffusion process, we should allocate the inference-time computation.  [^1] Website: https://sites.google.com/view/t2v-dlbs
Inference-Time Search, Text-to-Video Generation, Diffusion Models
[…]
tcb@breakable
@@ 1 Introduction
The remarkable progress in text-to-video diffusion models enables photorealistic, high-resolution video generation (OpenAI, 2024; Google DeepMind, 2024; Chen et al., 2024; Blattmann et al., 2023a). Many future applications are anticipated such as creating novel games (Bruce et al., 2024), movies (Zhu et al., 2023), or simulators to control real-world robots (Brooks et al., 2024). However, the detailed contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes, which should not happen in the real world. For instance, simulating factual physics in the generated video is still challenging (Liu et al., 2024; Bansal et al., 2024). Recently, it has attracted a lot of attention to steering the output of diffusion models based on reward evaluation quantifying the goodness of the content, which is studied as an alignment problem (Lee et al., 2023b; Huang et al., 2024a). There is a large room for improvement of perceptual quality along the frame direction in the video, and to align models with our preference, we should address which metrics we should optimize and how we can optimize them.
@@
Figure 1: Diffusion latent beam search (DLBS) seeks a better diffusion path over the reverse process; sampling $K$ latents per beam and possessing $B$ beams for the next step, which mitigates the effect from inaccurate $\operatorname{argmax}$. Lookahead estimator notably reduces the noise at latent reward evaluation by interpolating the rest of the time steps from the current latent with deterministic DDIM.
Figure 2: Qualitative evaluation among different search methods. We test prompts from DEVIL-high (above) and DEVIL-static (below). DLBS-LA generates more dynamic, and prompt-aligned videos than GS or BoN.
@@
In this paper, we propose Diffusion Latent Beam Search (DLBS) with lookahead estimator, an inference-time search over the reverse process (Figure 1), which can select better diffusion latent to maximize a given alignment reward. A lookahead estimator reduces the noise in reward estimate and a beam search robustly explores the latent paths avoiding inaccurate $\operatorname{argmax}$ operation.
@@
We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration of existing metrics (Huang et al., 2024b). When we evaluate outputs by leveraging capable vision language models (OpenAI, 2023; Gemini Team, 2023) as a proxy of human rater, many previous metrics to quantify the naturalness of video do not always correlate with each other, and optimal reward design to measure perceptual quality highly depends on the degree of dynamic descriptions in evaluation prompts. We design a weighted linear combination of multiple metrics, which is calibrated to perceptual quality and improves the correlation with preference.
@@
We demonstrate that our method can induce high-quality outputs based on the calibrated reward or AI feedback (Figure 2), without model parameter update, and outputs the best generation under the same search budget compared to greedy search (Kim et al., 2024; Huang et al., 2024a) and best-of-N sampling (Zhang et al., 2024b; Ma et al., 2025). We also provide practical efficiency guidelines on which axes, among search budget, lookahead steps for reward estimate, and denoising steps, in the denoising process we should allocate the inference-time computation.
@@ 2 Preliminaries
Latent Diffusion Models   Latent diffusion models (Rombach et al., 2022; Ma et al., 2024) are a special class of diffusion probabilistic models (Sohl-Dickstein et al., 2015; Ho et al., 2020), and popular choices for high-resolution text-to-video generation (He et al., 2022; Blattmann et al., 2023b, a), which considers the diffusion process in embedding space. Let $\textbf{x}_{0}$ as a video and encode it as $\textbf{z}_{0}=\text{Enc}(\textbf{x}_{0})$ using VAE (Kingma & Welling, 2013). Continuous-time forward diffusion process can be modeled as a solution to a stochastic differential equation (SDE) (Song et al., 2021b):
$d\textbf{z}=\textbf{f}(\textbf{z},t)dt+g(t)d\textbf{w},$ (1)
where $_{0}\sim p_{0}(\textbf{z})$ z ) is the latent as initial condition while $p_{t}(\textbf{z})$ is the marginal distribution of $\textbf{z}_{t}$, $\textbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g:\mathbb{R}\times\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\textbf{w}\in\mathbb{R}^{d}$ is $d$-dimensional standard Wiener process. $\textbf{f}(\cdot,\cdot)$ and $g(\cdot)$ are designed appropriately for the marginal distribution to reach $p_{T}(\textbf{z})\approx\mathcal{N}(0,\textbf{I})$ as $t\rightarrow T$ (Karras et al., 2022). Reverse diffusion process generates samples $\textbf{z}_{0}$ through the following reverse-time SDE:
$d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})]dt+g(t)d\bar{\textbf{w}},$ (2)
@@
where $dt$ here is an infinitesimal negative time step from $T$ to 00 and $\bar{\textbf{w}}\in\mathbb{R}^{d}$ is a standard reverse-time Wiener process. We start this with $\textbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$. This SDE induces the marginal distribution on the data $p^{\text{pre}}(\textbf{z})$ (i.e. pre-trained diffusion models). While we omit the notation for simplicity, we consider the text-to-video generation problem, where the diffusion process is conditioned on text prompts c.
@@
Alignment for Text-to-Video Diffusion Models In this paper, we define the alignment problem in text-to-video generation as increasing the probability of generating perceptually good video for humans, such as $\max\mathbb{E}[p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})]$ where $\mathcal{O}\in\{0,1\}$ represents if the generated video $\textbf{x}_{0}$ conditioned on c is perceptually higher quality or not. The common assumption is such a probability depends on a proxy scalar reward function $r(\textbf{x}_{0},\textbf{c})$ such as $p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})\propto\exp(\beta^{-1}r(\textbf{x}_{0},\textbf{c}))$ with $\beta\in\mathbb{R}$, and then the problem comes down to reward maximization. The proxy reward function may take generated video $\textbf{x}_{0}$ and a prompt c as inputs.
@@
Alignment as Stochastic Optimal Control   Previous works formulate such a reward maximization problem from the view of stochastic optimal control (Uehara et al., 2024; Huang et al., 2024a; Domingo-Enrich et al., 2024), where we aim to find an additional drift term $\textbf{u}(\cdot,\cdot)$ for the following reverse SDE:
$d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})+\textbf{u}(\textbf{z},t)]dt+g(t)d\bar{\textbf{w}}.$ (3)
For convenience, we adopt the change-of-variables as $\boldsymbol{\nu}_{t}:=\mathbf{z}_{T-t}$ and $\bar{\textbf{f}}(\boldsymbol{\nu},t):=\textbf{f}(\boldsymbol{\nu},t)-g(t)^{2}\nabla_{\boldsymbol{\nu}}\log p_{t}(\boldsymbol{\nu})$ because stochastic control is often based on the standard flow of time ($t:0\rightarrow T$), and then Equation 3 is written as:
$d\boldsymbol{\nu}=[\bar{\textbf{f}}(\boldsymbol{\nu},t)+\textbf{u}(\boldsymbol{\nu},t)]dt+g(t)d\textbf{w},$ (4)
where $dt$ here is an infinitesimal time step and $d\textbf{w}$ is a standard Wiener process.
@@
where $r^{\prime}(\cdot):=r(\text{Dec}(\cdot))$ evaluates the latent in the video space and $\lambda>0$. The expectation is taken over samples from Equation 4. In stochastic control, it is known that we can define the optimal value function,
@@
where $p^{*}_{t}(\boldsymbol{\nu})=\frac{1}{Z}\exp(\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda})p^{\text{pre}}_{t}(\boldsymbol{\nu})$, and obtain the optimal drift $\textbf{u}^{*}(\boldsymbol{\nu},t)=g(t)^{2}\nabla_{\boldsymbol{\nu}}\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda}$ (Pavon, 1989). This optimal value function is the solution of stochastic Hamilton-Jacobi-Bellman equation (Evans, 2022) according to this Feynman-Kac formula (Øksendal, 2003; Moral, 2004):
@@
and then we obtain a tractable form of optimal drift term as:
@@
The intuition here is the optimal drift pulls the current latent $\boldsymbol{\nu}$, while following the pre-trained reverse SDE, into the region achieving a higher reward at time $T$.

### v2

@@ Abstract
The remarkable progress in text-to-video diffusion models enables photorealistic generations, although the contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes. Recently, an alignment problem has attracted huge attention, where we steer the output of diffusion models based on some quantity on the goodness of the content. Because there is a large room for improvement of perceptual quality along the frame direction, we should address which metrics we should optimize and how we can optimize them in the video generation. In this paper, we propose diffusion latent beam search with lookahead estimator, which can select a better diffusion latent to maximize a given alignment reward, at inference time. We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration by weighting existing metrics. This is because when humans or vision language models evaluate outputs, many previous metrics to quantify the naturalness of video do not always correlate with evaluation. We demonstrate that our method improves the perceptual quality evaluated on the calibrated reward, VLMs, and human assessment, without model parameter update, and outputs the best generation compared to greedy search and best-of-N sampling under much more efficient computational cost. The experiments highlight that our method is beneficial to many capable generative models, and provide a practical guideline that we should prioritize the inference-time compute allocation into lookahead steps for reward estimation over search budget or denoising steps.  [^1] Website: https://sites.google.com/view/t2v-dlbs  [^2] Code: https://github.com/shim0114/T2V-Diffusion-Search
@@ 1 Introduction
The remarkable progress in text-to-video diffusion models enables photorealistic, high-resolution video generation [1, 2, 3, 4]. Many future applications are anticipated, such as creating novel games [5], movies [6], or simulators to control real-world robots [7]. However, the detailed contents of the generated video often include unnatural movement or deformation, reverse playback, and motionless scenes, which should not happen in the real world. For instance, simulating factual physics in the generated video is still challenging [8, 9]. Recently, it has attracted a lot of attention to steering the output of diffusion models based on reward evaluation, quantifying the goodness of the content, which is studied as an alignment problem [10, 11]. There is a large room for improvement of perceptual quality along the frame direction in the video, and to align models with our preferences, we should address which metrics to optimize and how to optimize them.
@@
Figure 1: (Left) Diffusion latent beam search (DLBS) seeks a better diffusion path over the reverse process; sampling $K$ latents per beam and possessing $B$ beams for the next step, which mitigates the effect from inaccurate $\operatorname{argmax}$. Lookahead (LA) estimator notably reduces the noise at latent reward evaluation by interpolating the rest of the time steps from the current latent with deterministic DDIM. (Right) DLBS achieves much better computational-efficiency than best-of-N (BoN), as achieving higher performance gains under the same execution time. LA estimator (DLBS-LA) could remarkably boost efficiency only with marginal overhead on top of DLBS.
@@
In this paper, we propose Diffusion Latent Beam Search (DLBS) with lookahead estimator, an inference-time search over the reverse process (Figure 1; Left), which can select a better diffusion latent to maximize a given alignment reward. A lookahead estimator reduces the noise in the reward estimate, and a beam search robustly explores the latent paths, avoiding inaccurate $\operatorname{argmax}$ operation.
@@
We then point out that the improvement of perceptual video quality considering the alignment to prompts requires reward calibration of existing metrics [12]. When we evaluate outputs by capable vision language models [13, 14] or human raters, many previous metrics to quantify the naturalness of video do not always correlate with each other, and optimal reward design to measure perceptual quality highly depends on the degree of dynamic descriptions in evaluation prompts. We design a weighted linear combination of multiple metrics, which is calibrated to perceptual quality and improves the correlation with preference.
@@
We demonstrate that DLBS can induce high-quality outputs based on the calibrated reward, AI and human feedback (Figure 2), without model parameter update, and realize the best generation under much more efficient computational cost compared to greedy search [15, 11] and best-of-N sampling [16, 17]. The experiments also highlight that our method is beneficial to many SoTA models (e.g., Latte [18], CogVideoX [19], and Wan 2.1 [20]), and provide a practical guideline on that we should prioritize the compute allocation into lookahead steps for reward estimate more than search budget or denoising steps, at an inference time.
@@ 2 Preliminaries
Latent Diffusion Models   Latent diffusion models [21, 18] are a special class of diffusion probabilistic models [22, 23], and popular choices for high-resolution text-to-video generation [24, 25, 4], which considers the diffusion process in embedding space. Let $\textbf{x}_{0}$ be a video and encode it as $\textbf{z}_{0}=\text{Enc}(\textbf{x}_{0})$ using VAE [26]. Continuous-time forward diffusion process can be modeled as a solution to a stochastic differential equation (SDE) [27]: $d\textbf{z}=\textbf{f}(\textbf{z},t)dt+g(t)d\textbf{w},$ where $_{0}\sim p_{0}(\textbf{z})$ z ) is the latent as initial condition while $p_{t}(\textbf{z})$ is the marginal distribution of $\textbf{z}_{t}$, $\textbf{f}:\mathbb{R}^{d}\times\mathbb{R}\rightarrow\mathbb{R}^{d}$ is the drift coefficient, $g:\mathbb{R}\rightarrow\mathbb{R}$ is the diffusion coefficient, and $\textbf{w}\in\mathbb{R}^{d}$ is $d$-dimensional standard Wiener process. $\textbf{f}(\cdot,\cdot)$ and $g(\cdot)$ are designed appropriately for the marginal distribution to reach $p_{T}(\textbf{z})\approx\mathcal{N}(0,\textbf{I})$ as $t\rightarrow T$ [28]. Reverse diffusion process generates samples $\textbf{z}_{0}$ through the following reverse-time SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})]dt+g(t)d\bar{\textbf{w}},$ where $dt$ here is an infinitesimal negative time step from $T$ to 00 and $\bar{\textbf{w}}\in\mathbb{R}^{d}$ is a standard reverse-time Wiener process. We start this with $\textbf{z}_{T}\sim\mathcal{N}(0,\textbf{I})$. This SDE induces the marginal distribution on the data $p^{\text{pre}}(\textbf{z})$ (i.e., pre-trained diffusion models). While we omit the notation for simplicity, we consider the text-to-video generation problem, where the diffusion process is conditioned on text prompts c.
@@
Alignment for Text-to-Video Diffusion Models In this paper, we define the alignment problem in text-to-video generation as increasing the probability of generating perceptually good video for humans, such as $\max\mathbb{E}[p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})]$ where $\mathcal{O}\in\{0,1\}$ represents if the generated video $\textbf{x}_{0}$ conditioned on c is perceptually higher quality or not. The common assumption is such a probability depends on a proxy scalar reward function $r(\textbf{x}_{0},\textbf{c})$ such as $p(\mathcal{O}=1|\textbf{x}_{0},\textbf{c})\propto\exp(\beta^{-1}r(\textbf{x}_{0},\textbf{c}))$ with $\beta\in\mathbb{R}$, and then the problem comes down to reward maximization. The proxy reward function may input the generated video $\textbf{x}_{0}$ and a prompt c.
@@
Alignment as Stochastic Optimal Control   Previous works formulate such a reward maximization problem from the view of stochastic optimal control [29, 11, 30], where we aim to find an additional drift term $\textbf{u}(\cdot,\cdot)$ for the following reverse SDE: $d\textbf{z}=[\textbf{f}(\textbf{z},t)-g(t)^{2}\nabla_{\textbf{z}}\log p_{t}(\textbf{z})+\textbf{u}(\textbf{z},t)]dt+g(t)d\bar{\textbf{w}}$. For convenience, we adopt the change-of-variables as $\boldsymbol{\nu}_{t}:=\mathbf{z}_{T-t}$ and $\bar{\textbf{f}}(\boldsymbol{\nu},t):=\textbf{f}(\boldsymbol{\nu},t)-g(t)^{2}\nabla_{\boldsymbol{\nu}}\log p_{t}(\boldsymbol{\nu})$ because stochastic control is often based on the standard flow of time ($t:0\rightarrow T$), and then the original SDE is re-written as: $d\boldsymbol{\nu}=[\bar{\textbf{f}}(\boldsymbol{\nu},t)+\textbf{u}(\boldsymbol{\nu},t)]dt+g(t)d\textbf{w}$, where $dt$ here is an infinitesimal time step and $d\textbf{w}$ is a standard Wiener process.
@@
where $r^{\prime}(\cdot):=r(\text{Dec}(\cdot))$ evaluates the latent in the video space and $\lambda>0$. $\mathbb{E}[\cdot]$ is taken over sampling process above. In stochastic control, the optimal value function is known to be defined as,
@@
where $p^{*}_{t}(\boldsymbol{\nu})=\frac{1}{Z}\exp(\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda})p^{\text{pre}}_{t}(\boldsymbol{\nu})$, and obtain the optimal drift $\textbf{u}^{*}(\boldsymbol{\nu},t)=g(t)^{2}\nabla_{\boldsymbol{\nu}}\frac{v^{*}_{t}(\boldsymbol{\nu})}{\lambda}$ [31]. This optimal value function is the solution of stochastic Hamilton-Jacobi-Bellman equation [32] according to this Feynman-Kac formula [33, 34]:
@@
and then we obtain a tractable form of the optimal drift term as:
@@
The intuition here is that the optimal drift pulls the current latent $\boldsymbol{\nu}$, while following the pre-trained reverse SDE, into the region achieving a higher reward at time $T$.
Figure 2: Qualitative evaluation among different search methods. We test prompts from DEVIL-high (above) and DEVIL-static (below). DLBS-LA generates more dynamic and prompt-aligned videos than GS or BoN.
3 Diffusion Latent Beam Search
We first provide a unified view of existing inference-time alignment methods through several practical approximations of optimal drift $\mathbf{u}^{*}(\boldsymbol{\nu}_{t},t)$ (Section 3.1). To mitigate errors from approximations, we propose a novel search algorithm, diffusion latent beam search with lookahead estimator (Section 3.2).
3.1 A Unified View on Practical Approximations
While Equation 4 has a relatively tractable form, it is still computationally expensive, since the expectation requires complete diffusion sampling to evaluate the latent at each time step and face numerical instability. Previous alignment methods rely on multiple-step practical approximations.
Step. 1: Jensen’s Inequality   First, when assuming $\frac{r^{\prime}(\boldsymbol{\nu}_{T})}{\lambda}$ is almost deterministic (this might hold when $t\rightarrow T$), Jensen’s inequality yields the following approximation by exchanging $\log$ and $\mathbb{E}[\cdot]$, which can be considered as a certain form of classifier guidance [35]:
@@ 3.1 A Unified View on Practical Approximations
Step. 2: Tweedie’s Formula   To avoid the computationally expensive expectation, the expected reward is further approximated as $\mathbb{E}_{p^{\text{pre}}}[r^{\prime}(\boldsymbol{\nu}_{T})|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]\approx r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ where $\hat{\boldsymbol{\nu}}_{T|t}\approx\mathbb{E}_{p^{\text{pre}}}[\boldsymbol{\nu}_{T}|\boldsymbol{\nu}_{t}=\boldsymbol{\nu}]$ is a one-step approximation of posterior mean [36], which can be calculated only with the current latent $\boldsymbol{\nu}_{t}$ without full diffusion path. Therefore, the optimal drift term can be seen as solely depending on the current time step $t$:
Such a computationally tractable drift term has been leveraged for previous inference-time alignment methods via approximate guidance or twisted sequential Monte Carlo (SMC) [37]. However, as the approximated posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$in intermediate steps is noisy, evaluation with the reward function for clean data $r^{\prime}(\cdot)$ may not provide a reliable signal [38]. Moreover, Equation 6 requires the reward gradient, which is not applicable to non-differentiable rewards, such as AI feedback, and is also not suitable for modalities whose reward gradient imposes a huge computational cost in practice, such as video.
Step. 3: Converting Reward Gradient into argmax   The usage of reward gradient can be converted into $\operatorname{argmax}$ operator [11, 39, 40]. The intuition here is that since the optimal drift in Equation 6 induces the diffusion latent to the direction where it maximizes the reward, we replace such a maximization with a zeroth-order search. The SDE is approximated as:
$d\boldsymbol{\nu}=\bar{\textbf{f}}(\boldsymbol{\nu},t)dt+g(t)d\textbf{w}^{*}~{}~{}\text{where}~{}~{}d\textbf{w}^{*}=\operatorname{argmax}_{d\textbf{w}}~{}r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t}).$ (7)
Note that the current diffusion latents $\boldsymbol{\nu}_{t}$ and posterior mean $\hat{\boldsymbol{\nu}}_{T|t}$ are sampled by following the standard Wiener process $d\textbf{w}$. This approximation is leveraged for inference-time alignment via greedy search [11, 39] or SMC [41] of diffusion latents. However, greedy search can result in sub-optimal generation affected by inaccurate reward estimate $r^{\prime}(\hat{\boldsymbol{\nu}}_{T|t})$ due to its noisy input. Moreover, it can be challenging to obtain an accurate density ratio term required in SMC for a high-dimensional domain such as video generation.
3.2 Mitigating Approximation Errors via Beam Search
Existing practical algorithms based on these three approximations, such as greedy search [11, 39], fall into sub-optimal generation due to the erroneous reward evaluation with a noisy estimate of the posterior mean [36], and $\operatorname{argmax}$ operator based on them. To resolve the error accumulation, we propose a simple yet robust modification, diffusion latent beam search (DLBS) with lookahead estimator. To clearly describe the practical implementation, we use the notation of a discrete-time diffusion process in the rest of the section (see Appendix D for the continuous-time diffusion process).
Practical Implementation   We summarize the detailed sampling procedure of DLBS in Algorithm 1. For the diffusion sampler, we use stochastic DDIM [42] with a decreasing sequence $\{\alpha_{t}\}_{t=1}^{T}\in(0,1]^{T}$, noise level $\eta$, and noise schedule $\sigma_{t}=\eta\sqrt{(1-\alpha_{t-1})/(1-\alpha_{t})}\sqrt{1-\alpha_{t-1}/\alpha_{t}}$, which is equivalent to DDPM [23] when $\eta=1.0$. We initialize $B$ latent beams from the Gaussian distribution (Line 1.1), sample $K$ latents per beam in the next time step (Line 1.11), and then compute the one-step estimation of the posterior mean (Line 1.13). DLBS evaluates the estimator of posterior mean $\hat{\mathbf{z}}_{0|t-1}$ with reward function (Line 1.18) and selects Top-$B$-rewarded latent beams instead of Top-$1$ (i.e., $\operatorname{argmax}$) from $KB$ candidates (Line 1.19), which is iterated over entire reverse process from $t=T$ to $t=0$. DLBS can possess latent beams more widely than greedy search under the same budget, which mitigates error propagation due to the approximated diffusion latent evaluation.